OpenStudy (anonymous):

(sin Θ − cos Θ)2 − (sin Θ + cos Θ)2

OpenStudy (anonymous):

−4sin(Θ)cos(Θ) 2 sin2 Θ cos2 Θ

OpenStudy (dumbcow):

distribute, then combine terms note : sin^2 + cos^2 = 1

OpenStudy (unklerhaukus):

use difference of squares, a^2 - b^2 = (a+b)(a-b)

OpenStudy (midhun.madhu1987):

\[a ^{2} - b ^{2} = (a-b)(a+b)\]

OpenStudy (unklerhaukus):

where a = sin Θ − cos Θ b = sin Θ + cos Θ

OpenStudy (unklerhaukus):

can you do it @gabbimanges17 ?

OpenStudy (anonymous):

@UnkleRhaukus is the answer sin^2?

OpenStudy (unklerhaukus):

nope. what do you get for (a-b)? and what do you get for (a+b)?

OpenStudy (anonymous):

@UnkleRhaukus im not sure

OpenStudy (unklerhaukus):

\[ (\sin\theta − \cos\theta)^2 − (\sin\theta + \cos\theta)^2\] \[\Big((\sin\theta − \cos\theta)+(\sin\theta + \cos\theta)\Big)\Big((\sin\theta − \cos\theta)-(\sin\theta + \cos\theta)\Big)\]

OpenStudy (anonymous):

looks like they cancel each other out

OpenStudy (unklerhaukus):

many of the terms cancel away, but not all of them

OpenStudy (anonymous):

@UnkleRhaukus can you specify?

OpenStudy (unklerhaukus):

\[\color{teal}{(\sin\theta − \cos\theta)}^2 − \color{orange}{(\sin\theta + \cos\theta)}^2\] \[=\Big(\color{teal}{(\sin\theta − \cos\theta)}+\color{orange}{(\sin\theta + \cos\theta)}\Big)\Big(\color{teal}{(\sin\theta − \cos\theta)}-\color{orange}{(\sin\theta + \cos\theta)}\Big)\\ =\Big(\sin\theta − \cos\theta+\sin\theta + \cos\theta\Big)\Big(\sin\theta − \cos\theta-\sin\theta - \cos\theta\Big)\\ = \]

OpenStudy (unklerhaukus):

\[=\Big(\sin\theta +\sin\theta + \cos\theta − \cos\theta\Big)\Big(\sin\theta -\sin\theta - \cos\theta− \cos\theta\Big)\\ =\]

OpenStudy (anonymous):

0

OpenStudy (unklerhaukus):

work it out properly

OpenStudy (anonymous):

im not sure what youre asking.. it looks like the last two cos dont cancel out

OpenStudy (unklerhaukus):

\[=\Big(\sin\theta +\sin\theta + \cos\theta − \cos\theta\Big)\Big(\sin\theta -\sin\theta - \cos\theta− \cos\theta\Big)\\ =\Big((1+1)\sin\theta + (1-1)\cos\theta\Big)\Big((1-1)\sin\theta +(-1-1)\cos\theta\Big)\\ =\]

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

you have me in a twist

OpenStudy (unklerhaukus):

simplify!

OpenStudy (anonymous):

−4sin(Θ)cos(Θ)

OpenStudy (unklerhaukus):

are you sure?

OpenStudy (anonymous):

im not sure about anything anymore, please can you explain in plain english

OpenStudy (unklerhaukus):

The expression given is a difference of squares, we used a formula we already knew for the difference of square to rearrange the expression. Then we simplified the terms in this new form, and we got ...

OpenStudy (anonymous):

−4sin(Θ)cos(Θ)

OpenStudy (unklerhaukus):

good.

OpenStudy (anonymous):

thank you!